common factor, set equal to o, then solve.
2x^2+6x-56=0

Answer by stanbon(57346)

common factor, set equal to o, then solve.
2x^2+6x-56=0
----
2(x^2+3x-28) = 0
---
x^2+3x-28 = 0
(x+7)(x-4) = 0
x = -7 or x = 4
=====================
Cheers,
Stan H.
=====================